Please help me solve this problem. New cars depreciate an average of 28% in thier first year. What will a $9000 car be worth after 1 year?: Please help me solve this problem. New cars depreciate an average of 28% in thier first year. What will a $9000 car be worth after 1 year?
Answer by checkley77(3848)

9,000-.28*9,0000=9,000-2,520=$6,480 is the depreciated value of the car affter 1 year.